How much of an alloy that is 10% copper should be mixed with 600 oz of an alloy that is 60% copper in order to get an alloy that is 20% cooper
Answer by mananth(16946)

--- percent ---------------- Amount
Alloy I 60 ---------------- 600 oz
Alloy II 10 ---------------- x oz
Mixture 20 ---------------- 600 + x oz
60% * 600 + 10% x = 20% ( 600 + x )
multiply by 100

60 * 600 + 10 x = 20 ( 600 + x )
36000 + 10 x = 12000 + 20 x
10 x -20 x = -36000 + 12000
-10 x = -24000
/ -10
x = 2400 oz Alloy II

m.ananth@hotmail.ca